OpenStudy (anonymous):

HELPPPPPP! Find the exact value of sin 7pi/4

OpenStudy (nathan917):

sin(7pi/4) = sin(2pi -- pi/4) = sin(-pi/4) = -sin(pi/4) = - sqrt(2)/2

OpenStudy (anonymous):

If you take a look at the unit circle I attached, it will help you with future problems as well. I circled sin (7pi/4)